### v1

# FMNet: Frequency-Assisted Mamba-Like Linear Attention Network for Camouflaged Object Detection

Abstract
Camouflaged Object Detection (COD) is challenging due to the strong similarity between camouflaged objects and their surroundings, which complicates identification. Existing methods mainly rely on spatial local features, failing to capture global information, while Transformers increase computational costs.To address this, the Frequency-Assisted Mamba-Like Linear Attention Network (FMNet) is proposed, which leverages frequency-domain learning to efficiently capture global features and mitigate ambiguity between objects and the background. FMNet introduces the Multi-Scale Frequency-Assisted Mamba-Like Linear Attention (MFM) module, integrating frequency and spatial features through a multi-scale structure to handle scale variations while reducing computational complexity. Additionally, the Pyramidal Frequency Attention Extraction (PFAE) module and the Frequency Reverse Decoder (FRD) enhance semantics and reconstruct features. Experimental results demonstrate that FMNet outperforms existing methods on multiple COD datasets, showcasing its advantages in both performance and efficiency. Code available at https://anonymous.4open.science/r/FMNet-3CE5.
Index Terms:
camouflaged object detection, semantic enhancement, mamba-like linear attention, frequency-assisted
I Introduction
Camouflaged Object Detection (COD) is a highly challenging task aimed at accurately detecting hidden objects that closely resemble their surrounding environment. COD has significant application value in various fields, such as lesion segmentation in medical imaging and defect detection in industrial settings.
Traditional COD methods primarily rely on handcrafted feature extraction[1][2]. While these methods achieve certain performance in specific scenarios, their robustness is limited. With the availability of open-source COD datasets[3][4][5], deep learning-based COD methods have shown significant advantages by automatically extracting rich features[3][6]. Recent studies have demonstrated that optimizing feature representation can effectively address challenges such as varying target sizes, occlusion by surroundings, and blurred boundaries in camouflaged objects. Strategies such as multi-scale feature extraction[7][8] and boundary guidance[9] have been proposed. However, most existing methods are still constrained to optimizing local features, making it difficult to capture global information effectively. This limitation is particularly pronounced when detecting camouflaged objects with significant size variations or occlusions. Additionally, spatial domain features are prone to interference from complex backgrounds, often caused by overemphasizing local details or individual pixel positions.
Figure 1: Comparison of the proposed method with others using traditional attention mechanisms.
Figure 2: Overview of the proposed FMNet framework, proposed method adopts MambaVision as the encoder and introduces the following core modules: the Pyramidally Frequency Attention Extraction (PFAE) module based on frequency domain attention, the Multi-scale Frequency-Assisted Mamba-Like Linear Attention (MFM) module, and the Frequency Reverse Decoder (FRD).
Transformer methods [10], with their ability to model long-range dependencies, have been widely applied in camouflaged object detection. However, the high computational cost and complex network structure of Transformers [11][12] significantly limit their practicality.
Frequency-domain features [13][14] have gained significant attention in camouflaged object detection due to their inherent global modeling capabilities, which effectively suppress background noise and improve the semantic clarity of camouflaged objects. This advantage is particularly evident in scenarios with blurred boundaries and occlusions, where camouflaged objects are difficult to distinguish. However, frequent transformations between frequency and spatial domains lead to increased computational complexity and parameter overhead.
In recent years, Mamba methods [15][16], with their efficient attention mechanisms and lightweight design, have significantly reduced computational costs and shown great application potential. However, the potential of Mamba methods in camouflaged object detection remains underexplored.
Based on the above discussion, a novel method named Frequency-Assisted Mamba-Like Linear Attention Network (FMNet) is proposed. This method integrates feature information from both the frequency and spatial domains, introduces a multi-scale strategy to further extract global information, and leverages Mamba-Like Linear Attention (MLLA)[17] to optimize popular Transformer-based methods for COD. The comparison between our method and Transformer-based COD methods is illustrated in Fig. 1. The main contributions in this paper can be summarized as follows.
  • FMNet is proposed, distinct from traditional Transformer methods, by incorporating multi-scale and frequency-assisted strategies to significantly enhance the performance of camouflaged object detection.
  • The Multi-scale Frequency-Assisted Mamba-Like Linear Attention (MFM) module is designed, which synergizes frequency-domain and spatial features, providing a more comprehensive understanding of image characteristics while effectively reducing computational complexity.
  • The Pyramidal Frequency Attention Extraction (PFAE) module and the Frequency Reverse Decoder (FRD) module are innovatively developed to enhance the representation of frequency-domain features and integrate multi-level information, further improving detection performance.
II Methodology
II-A Overview
The proposed FMNet is shown in Fig. 2. Given an input image $I_{c}\in\mathbb{R}^{H\times W\times 3}$, a hybrid backbone integrating Transformer and Mamba[18] is adopted to efficiently extract the initial features $E_{i}$, where the resolution of each feature map is progressively reduced to $\frac{1}{2^{i+1}}$ of its original size. Inspired by previous works [14][19], a PFAE is employed to extract multi-scale fused features $E_{5}$. To efficiently model long-range dependencies across the feature domain, the MFM is used to refine the global context and produce optimized features $F_{i}$. Finally, FRD aggregates these multi-level features to produce the final feature maps $G_{i}$.
Figure 3: Details of the Multi-scale Frequency-Assisted Mamba-Like Linear Attention(MFM).
Figure 4: Details of Frequency Reverse Decoder.
II-B Pyramidally Frequency Attention Extraction
The proposed PFAE module, as shown in Fig. 2, integrates attention mechanisms in the frequency domain, allowing for more effective extraction of high-frequency features.
Specifically, the input feature $E_{4}$ is first passed through a $1\times 1$ convolution to reduce the number of channels, resulting in $\hat{E_{4}}$. Subsequently, four dilated convolution layers with dilation rates $z$ are loaded as four branches:
$\tilde{I_{n}}=C_{1}AC_{z}(\hat{E_{4}}+I_{n-1}),\quad z=2n-1,\quad n\geq 2$ (1)
each containing a frequency-domain attention module, this module uses the Fast Fourier Transform (FFT) to generate queries $Q$, keys $K$, and values $V$. After reshaping, the queries and keys are dot-multiplied to generate the transpose attention map $A^{\prime}_{f}$:
$Q,K,V=fft(\hat{E_{4}}),\quad A^{\prime}_{f}=\tilde{Q}\odot\tilde{K}$ (2)
where $\hat{Q}$ and $\hat{K}$ are the results of applying the reshape operation to the query $Q$ and key $K$, since $A_{f}$ is of complex type, activate its real and imaginary parts separately, and then merge them:
$A_{f}^{re}=\frac{A^{\prime}_{f}+\text{coj}(A^{\prime}_{f})}{2},\quad A_{f}^{im}=\frac{A^{\prime}_{f}-\text{coj}(A^{\prime}_{f})}{2i}$ (3)
$A_{f}=\Theta(\text{Sof}(A_{f}^{re}),\text{Sof}(A_{f}^{im}))$ (4)
where $\Theta(\cdot,\cdot)$ represents a function that combines the real and imaginary parts into a complex number, and $\text{Sof}(\cdot)$ represents the Softmax function. Subsequently, the attention map $A_{f}$ performs a dot product with $V$ to achieve weighted optimization, and the result is converted back to the original domain via the Inverse Fast Fourier Transform (IFFT). During this process, a Frequency Weight Module(FWM) is introduced, which will be discussed in Section II-C, to perform residual connections. Finally, a $1\times 1$ convolution is applied, and a residual connection with the features before the FFT transformation is added to generate the hybrid feature $I_{n}$:
$I_{n}=C_{1}\Phi(\|\text{ifft}(A_{f}\odot V)\|,FWM)+\tilde{I_{n}}$ (5)
where $\Phi$ represents the concatenation operation. The output $E_{5}$ is generated through convolution operations:
$E_{5}=C_{3}C_{1}(\Phi(J_{1},J_{2},J_{3},J_{4})+P_{4}^{128})$ (6)
where $C_{k}$ denotes a convolution operation with a kernel size of $k\times k$.
II-C Multi-scale Frequency-Assisted Mamba-Like Linear Attention
The multi-head attention module[10] decomposes input queries, keys, and values into $N_{h}$ parts along the channel dimension, self-attention learning is performed based on the queries, keys, and values using the formula:
$A=\sigma\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{D_{h}}}\right)V_{i}$ (7)
where $\sigma$ denotes the Sigmoid function, it can be seen that for an input feature map $X\in\mathbb{R}^{H\times W\times D}$, the computational complexity is $\mathcal{O}(H^{2}W^{2}D)$. Evidently, this complexity increases rapidly as the image resolution increases. As a result, many camouflaged object detection networks that adopt similar modules inevitably face high computational costs[12][20]. Inspired by methods aiming to address this issue[21][15], a module named Multi-scale Frequency-Assisted Mamba-Like Linear Attention (MFM) is proposed.
Linearized attention. Linear attention[21] replaces the non-linear Softmax function with linear normalization and introduces an additional kernel function $\phi$ into the query $q_{j}$ and key $k_{k}$. The output $y_{j}$ can be implemented as an autoregressive model, which is expressed as follows:
$S_{j}=S_{j-1}+K_{j}^{\top}V_{j},\quad Z_{j}=Z_{j-1}+K_{j}^{\top},\quad y_{j}=\frac{Q_{j}S_{j}}{Q_{j}Z_{j}}$ (8)
TABLE I: Comparison with State-of-the-Art Methods on three COD datasets. The symbols “↑”/“↓” indicate that higher/lower values are better. The best results are highlighted in red, and the second-best results are highlighted in blue.
Model Pub./Year CAMO-test COD10K-test NC4K
$S_{m}\uparrow$ $F_{\varphi}\uparrow$ $E_{m}\uparrow$ $M\downarrow$ $S_{m}\uparrow$ $F_{\varphi}\uparrow$ $E_{m}\uparrow$ $M\downarrow$ $S_{m}\uparrow$ $F_{\varphi}\uparrow$ $E_{m}\uparrow$ $M\downarrow$
JSOCOD[22] CVPR21 0.800 0.779 0.872 0.023 0.807 0.705 0.882 0.035 0.841 0.803 0.966 0.047
UGTR[23] ICCV21 0.785 0.738 0.823 0.086 0.818 0.712 0.853 0.035 0.839 0.787 0.874 0.052
ZoomNet[8] CVPR22 0.820 0.794 0.877 0.066 0.838 0.766 0.888 0.029 0.853 0.818 0.896 0.048
SINet-V2[7] TPAMI22 0.820 0.782 0.882 0.070 0.815 0.718 0.887 0.037 0.847 0.805 0.903 0.048
FRINet[24] ACMMM23 0.865 0.848 0.924 0.046 0.864 0.810 0.930 0.023 0.889 0.866 0.937 0.030
FSPNet[12] CVPR23 0.857 0.830 0.899 0.050 0.851 0.769 0.895 0.026 0.879 0.843 0.915 0.035
VSCode[25] CVPR24 0.873 0.844 0.925 0.046 0.869 0.806 0.931 0.023 0.891 0.863 0.935 0.032
GLCONet[11] TNNLS24 0.880 0.864 0.940 0.038 0.860 0.784 0.929 0.023 0.886 0.858 0.942 0.030
Ours - 0.890 0.863 0.946 0.039 0.895 0.798 0.940 0.020 0.890 0.871 0.944 0.029
TABLE II: Ablation analysis of the proposed method, with the best results highlighted in red.
No. Baseline PFAE MFM FRD Params (M) FLOPs (G) COD10K-Test NC4K
$S_{m}\uparrow$ $F_{\varphi}\uparrow$ $E_{m}\uparrow$ $M\downarrow$ $S_{m}\uparrow$ $F_{\varphi}\uparrow$ $E_{m}\uparrow$ $M\downarrow$
(a) 50.14 29.11 0.823 0.702 0.894 0.035 0.855 0.801 0.913 0.044
(b) 50.97 30.37 0.831 0.713 0.892 0.033 0.860 0.821 0.914 0.043
(c) 62.10 41.45 0.881 0.756 0.922 0.025 0.887 0.846 0.930 0.034
(d) 53.01 31.76 0.865 0.742 0.907 0.030 0.875 0.842 0.919 0.038
(e) 62.93 42.75 0.883 0.779 0.919 0.024 0.885 0.852 0.929 0.032
(f) 56.93 33.05 0.873 0.755 0.913 0.028 0.877 0.833 0.923 0.036
(g) 68.06 44.13 0.893 0.787 0.936 0.022 0.883 0.865 0.942 0.030
(h) 68.89 45.39 0.895 0.798 0.940 0.020 0.890 0.871 0.944 0.029
Selective state space model. The classical state space model maps the input $x(t)\in\mathbb{R}$ to the output $y(t)\in\mathbb{R}$ through the hidden state $h(t)\in\mathbb{R}^{d\times 1}$. Mamba[15] introduces the selective state space model, which discretizes the parameters using zero-order hold (ZOH). MLLA[17] rewrites Mamba into the following equivalent form:
$h_{j}=\widetilde{A}_{j}\odot h_{j-1}+B_{j}(\Delta_{j}\odot x_{j}),\quad y_{j}=C_{j}h_{j}+D\odot x_{j}$ (9)
where $\odot$ denotes the element-wise multiplication, $B_{j},C_{j},\Delta_{j}$ are derived from the input, $x_{j},\Delta_{j}\in\mathbb{R}^{1\times C}$, and $y_{j}\in\mathbb{R}^{1\times C}$.
It can be observed that (8) and (9) have a close relationship, specifically: $h_{j}\sim S_{j}\in\mathbb{R}^{d\times C},B_{j}\sim K_{j}^{\top}\in\mathbb{R}^{d\times 1},x_{j}\sim V_{j}\in\mathbb{R}^{1\times d},C_{j}\sim Q_{j}\in\mathbb{R}^{1\times d}$, $\tilde{A}_{i}$ plays the role of a forget gate in the Selective State Space Model (SSM). The MLLA module inherits the advantages of Mamba, adopting a structure similar to Mamba and introducing the forget gate mechanism. At the same time, to better adapt to vision tasks, MLLA replaces the traditional forget gate with positional encodings LePE[26], RoPE[27] and CPE[28]:
$LePE(x)=x+DWConv(x)W_{L}$ (10)
$RoPE(x_{m},\theta_{i})=x_{m}\cdot(\cos(m\theta_{i})+\sin(m\theta_{i}))$ (11)
$L_{a}=Att(RoPE(Q),RoPE(K),V+LePE(V))$ (12)
where $W_{L}$ is a learnable weight matrix, $DWConv$ denotes a depth-wise convolution, $x_{m}$ is the $m$-th dimension of the input, and $\theta_{i}$ is the angle related to the position, the complexity of the Linear Attention module is $O(NCd)=O(HWDD_{h})$.
Frequency-assisted Multi-scale MLLA. The proposed MFM module is shown in Fig. 3, where the figure omits reshaping operations for clarity. The optimization of MLLA allows us to design a multi-scale structure similar to [29] with acceptable complexity. The input feature $E_{i}$ is first processed through CPE and layer normalization to obtain the tensor $\tilde{E_{i}}$. Then, it is split along the channel dimension and processed through $1\times 1$ convolutions, followed by $n\times n$ depth-wise convolutions to generate tensors of different scales. These tensors are concatenated together after passing through the Linear Attention layer. The overall process is represented by the following formula:
$A_{i}^{n}=L_{a}(R(\sigma(D_{n}C_{1}(\tilde{E_{i}})))$ (13)
$A_{i}=\Phi(A_{i}^{3},A_{i}^{5})\quad F_{i}^{1}=L(A_{i}\odot R(\sigma(C_{1}(\tilde{E_{i}})))$ (14)
where $R$ represents the reshape operation, $D_{n}$ denotes the depthwise convolution kernel, $C_{1}$ refers to the convolution kernel with a size of 1, $\Phi$ represents the concatenation operation, and $F_{i}^{1}$ is the feature map output from the first stage.
A Frequency Weight Module (FWM) to enhance the representation of frequency-domain information through frequency residual connections is also designed:
$FWM=\|ifft(W(fft(X))*fft(X))\|$ (15)
where $W(\cdot)$ represents a series of operations consisting of convolution, batch normalization, GELU, convolution, and sigmoid functions applied sequentially, after passing through the multi-scale MLLA module, $F_{i}^{2}$ is obtained:
$F_{i}^{2}=CPE(F_{i}^{1}+FWM(E_{i})+E_{i})$ (16)
Finally, the overall output $F_{i}$ of the MFM can be obtained using the following formula:
$F_{i}=F_{i}^{2}+LN(Mlp(F_{i}^{2}))+FWM(F_{i}^{2})$ (17)
where $LN$ refers to Layer Normalization.
Figure 5: Visual comparison of the prediction maps with state-of-the-art methods, including large objects, small objects, and occluded objects.
II-D Frequency Reverse Decoder
The FRD module is shown in Fig. 4 Unlike existing COD methods [6] [12].
The input of the FRD consists of two parts: the auxiliary input and the main input $F_{i}$. The high-level feature map $G_{i+1}(F_{5})$ is used as an auxiliary input. First, bilinear interpolation and channel dimension expansion are applied to ensure that the auxiliary input matches the size and number of channels of the main input. Subsequently, the auxiliary input is concatenated with the main input along the channel dimension to obtain the feature $G_{i}^{1}$:
$G_{i}^{1}=\Phi(F_{i},Ex(G_{i+1}),...,Ex(Z)),\quad Z=\{G_{4},F_{5}\}$ (18)
the reverse attention is applied to the auxiliary features to generate a frequency-spatial hybrid reverse attention map $RA$, which is then used to produce the frequency-optimized reverse feature $G_{i}^{2}$:
$RA=\sum{(1-\sigma(G_{i+1}))+(1-\sigma(\|fft(G_{i+1})\|))}$ (19)
$G_{i}^{2}=RA*F_{i}$ (20)
finally, $G_{i}^{1}$ undergoes a series of convolution operations and is then integrated with $G_{i}^{2}$ to generate the final feature $G_{i}$, $G_{i}$ will serve as the auxiliary input for the next layer’s FRD.
$G_{i}=C_{3}\Phi(G_{i}^{1},Con(G_{i}^{2}))+F_{i}$ (21)
II-E Loss Function
Similar to these methods[6][12], the weighted binary cross-entropy (BCE) and weighted intersection over union (IoU)[30] are used as loss functions to supervise the multi-level features $G_{i}$. The loss function is defined as follows:
$\mathcal{L}_{\text{all}}=\sum_{i=1}^{5}2^{1-i}\left(\mathcal{L}_{\text{bce}}^{w}(G_{i},GT)+\mathcal{L}_{\text{iou}}^{w}(G_{i},GT)\right)$ (22)
where $\mathcal{L}_{\text{bce}}^{w}$ and $\mathcal{L}_{\text{iou}}^{w}$ represent the weighted BCE and IoU loss functions, respectively.
TABLE III: Efficiency analysis of the proposed FMNet and existing COD methods, P represents the parameters and F represents the FLOPs.
JSOCOD[22] FSPNet[12] GLCONet [11] Ours
P (M) 217.98 273.79 91.35 68.89
F (G) 112.34 283.31 173.58 45.39
III Experiments
III-A Experiment Setting
The FMNet method is evaluated on the CAMO[4], COD10K[3], and NC4K[5] datasets, using 4,000 images from the CAMO[4] and COD10K[3] dataset for training. The FMNet method is trained on three NVIDIA GTX 4090 GPUs (24GB) and utilizes the pre-trained MambaVision as an encoder to extract initial features. Input images are resized to $416\times 416$ and processed with data augmentation techniques such as random horizontal flipping and cropping. During training, the batch size is set to 30, over 100 epochs. The initial learning rate is set to $1e^{-4}$, and the Adam optimizer is used, with the learning rate reduced by a factor of 10 every 50 epochs. To evaluate method performance, the following metrics are adopted: S-measure ($S_{m}$), E-measure ($E_{m}$), Average F-measure ($F_{\varphi}$), and Mean Absolute Error ($\mathcal{M}$).
III-B Comparisons with the State-of-the-Art Methods
The FMNet is compared with 8 state-of-the-art methods, including JSOCOD[22], UGTR[23], ZoomNet[8], SINet-V2[7], FRINet[24], FSPNet[12], VSCode[25] and GLCONet[11]. All prediction results in this paper were provided by the original authors or obtained from open-source codes.
The quantitative comparison between the FMNet method and other SOTA methods is summarized in Tab. I. The results show that the method outperforms other methods in various evaluation metrics, demonstrating exceptional performance. Additionally, the parameters and FLOPs comparison between FMNet and other methods is provided in Tab. III. It can be seen that the proposed method demonstrates an efficiency advantage compared to complex networks.
Visual comparisons between FMNet and other methods are presented in Fig. 5, including segmentation results for large objects, small objects, and occluded objects. The results indicate that the proposed method demonstrates robust performance in all scenarios.
III-C Ablation Study
The quantitative results of the components are presented in the FMNet method in Table II. Specifically, the effectiveness of PFAE, MFM, and FRD is validated, as well as their impact on parameters and FLOPs. The experimental results show that adding the proposed modules to the baseline leads to a significant improvement in prediction performance. From Tab. II(e), Tab. II(f), and Tab. II(g), it can be observed that these three components exhibit good compatibility with each other.
Additionally, the efficiency ablation study reveals that the MLLA module consumes the most parameters and FLOPs while contributing the most to performance improvement, further proving that it is the most critical part for optimization.
IV Conclusion
This paper proposes a novel FMNet method for camouflaged object detection, with its core centered on a high-performance multi-scale frequency-assisted module (MFM). FMNet integrates feature information from both the spatial and frequency domains, achieving more precise object segmentation. Furthermore, a Pyramidally Frequency Attention Extraction (PFAE) module is designed to extract multi-scale features, and a Frequency Reverse Decoder (FRD) to cross-layer aggregation and reverse optimization, further improving the model performance. Extensive comparative experiments demonstrate that FMNet outperforms existing SOTA methods on multiple benchmark datasets. In the future, we will improve FMNet’s adaptability to complex environments and optimize it for edge device deployment to enhance its practical applications.